compound fard what is he empaival formula the eempoundm bills molar mass is 294.18gaool. What is the usoleeular formula of the conwnoundr

Please solve the whole question a and b. both of them are one question.
Answer #1

The compound has

26.58% Potassium

35.35% Chromium

38.07% Oxygen

a) For the imperical formula of the compound be KxCryOz

The atomic weight of potassium is 39

Chromium is 52

Oxygen is 16

molecular weight of the compound is (39x + 52y + 16z)

for potassium being 26.58% = 39x / (39x + 52y + 16z)

x = 6.81*10-3 *  (39x + 52y + 16z)

for chromium being 35.35% = 52y /  (39x + 52y + 16z)

y = 6.8*10-3 *  (39x + 52y + 16z)

for oxygen being 38.07% = 16z /  (39x + 52y + 16z)

z = 23.8*10-3 *  (39x + 52y + 16z)

Now by taking the ratio we get

x:y:z = [6.81*10-3 *  (39x + 52y + 16z)] :[ 6.8*10-3 *  (39x + 52y + 16z)] : [23.8*10-3 *  (39x + 52y + 16z)]

which can be reduced to

x:y:z = 1:1: 3.5

Hence the imperical formula for the given compound is  K1Cr1O3.5

b)

the molar mass of the compound is 294.18 g/mol

For the molecular formula of the compound be KaCrbOc

The atomic weight of potassium is 39

Chromium is 52

Oxygen is 16

molecular weight of the compound is (39a + 52b + 16c) = 294.18 g/mol

for potassium being 26.58% = 39a / (39a + 52b + 16c) = 39a/294.18 = 0.13a

a = 2

for chromium being 35.35% = 52b / (39a + 52b + 16c) = 0.18 b

b = 2

for oxygen being 38.07% = 16c / (39a + 52b + 16c) = 0.054c

c = 7

Hence the molecular formula for the given compound is  K2Cr2O7